# Stability Analysis of Cosmological Perturbations in the Bumblebee Model: Parameter Constraints and Gravitational Waves

### v2

@@ Abstract
We constrain the parameter space of the Bumblebee model within a cosmological background and investigate the properties of gravitational waves under the constrained parameter space. Specifically, we derive the conditions for the absence of ghost instability, Laplacian instability, and tachyonic instability for perturbations in a cosmological background. By incorporating the observed accelerated expansion of the universe and the observational constraints on tensor gravitational waves, we derive bounds on the parameter space of the Bumblebee model. We then examine the polarization modes, the propagation speeds, and the amplitude relations of gravitational waves within this constrained framework. Our results indicate that the non-minimal coupling parameter $\xi$ must be non-positive, and the Lorentz-violating parameter $\xi b^{2}$ has a lower bound on the order of $10^{-15}$. The propagation modes of gravitational waves in the Bumblebee model consist of two tensor modes, two vector modes, and a combination of two scalar modes. Notably, the tensor modes travel at subluminal speeds, whereas the vector and scalar modes propagate at superluminal speeds. These results provide a concrete theoretical framework and specific observational signatures for testing Lorentz invariance in the gravitational sector with future gravitational-wave detectors.
@@ I Introduction
Since its proposal, general relativity (GR) has undergone rigorous tests in both weak-field and strong-field regimes. In the weak-field limit, its predictions have been precisely verified through observations and experiments, such as the precession of Mercury’s perihelion einstein1915erklarung , the deflection of light dyson1923determination , and the Pound-Rebka experiment PhysRevLett.3.439 ; Pound:1960zz . In the strong-field limit, the orbital decay of the Hulse-Taylor pulsar shows excellent agreement with the gravitational wave (GW) radiation predicted by GR Hulse:1974eb ; Taylor:1979zz . The first direct detection of GWs, GW150914, confirms a key prediction of GR LIGOScientific:2016aoc ; LIGOScientific:2016emj . Imaging of the black hole shadows of M87* and Sagittarius A* by the Event Horizon Telescope provides further support for the black hole solutions predicted by GR EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022xqj . This series of observations and experiments has tested the validity of GR across both the weak-field regimes and the highly non-linear strong-field regimes.
@@
Although GR has successfully withstood many experimental tests, several critical issues remain difficult to explain within its framework, such as the dark matter problem Smith:1936mlg ; Zwicky:1937zza , the dark energy problem Peebles:2002gy , the problem of quantization tHooft:1974toh ; Goroff:1985th , and the hierarchy problem Arkani-Hamed:1998jmv ; Randall:1999ee ; Randall:1999vf . This has led to questions within the physics community about whether GR is the ultimate theory of gravity, which has spurred research into modified gravity theories.
@@
The main approaches for constructing modified gravity theories include Nojiri:2010wj ; Clifton:2011jh :
@@
Many of these modified gravity theories offer unique interpretations of GWs and the universe that differ from the interpretations provided by GR. For example, in general metric theories, up to six independent polarization modes of GWs are possible, significantly exceeding the two tensor modes predicted by GR Eardley:1973zuo . Certain modified gravity theories can effectively describe the evolution of the universe and provide a theoretical basis for explaining dark energy Nojiri:2006ri ; Nojiri:2017ncd . Some modified gravity theories fit galaxy rotation curves well, offering viable alternatives to the dark matter hypothesis Famaey:2011kh ; Capozziello:2012ie . Therefore, theoretical consistency analyses and experimental tests are crucial for identifying a more complete description of gravity.
@@
A theoretically consistent modified gravity theory must be stable. A critical issue in this regard is the ghost problem. If ghost degrees of freedom, i.e., field modes carrying negative energy, are present, the theory will inevitably lead to vacuum instability or even catastrophic decay Woodard:2006nt . Thus, the ghost-free condition serves as an important tool for selecting modified gravity theories and constraining theoretical parameters DeFelice:2016yws ; Kase:2018aps ; Clough:2022ygm . Especially, the ghost problem is almost ubiquitous in higher-order modified gravity theories. The well-known Ostrogradsky theorem Ostrogradsky:1850fid states that when a Lagrangian contains higher-order (second order or higher) time derivatives of dynamical variables, its Hamiltonian is usually bounded neither from above nor from below Woodard:2006nt . This implies that positive and negative parts could be excited to arbitrarily high energies, leading to vacuum instability Woodard:2006nt . There are also other studies concerning theoretical consistency, such as the initial value problem and the stability of evolution kreiss1989initial ; Papallo:2017qvl ; Bluhm:2008yt ; Clough:2022ygm .
@@
For the experimental tests of modified gravity theories, one of the most direct and promising approaches is the detection and characterization of GW signals. The polarization modes and propagation speeds of GWs are two critically important characteristics that typically vary across different modified gravity theories, making them direct probes for testing gravity theories. In GR, GWs propagate at the speed of light and exhibit only two tensor polarization modes: the plus mode ($P_{+}$) and the cross mode ($P_{\times}$). However, a general metric theory in four-dimensional spacetime allows for GWs up to six independent polarization modes (the additional ones being breathing ($P_{b}$), longitudinal ($P_{l}$), vector-$x$ ($P_{x}$), and vector-$y$ ($P_{y}$) modes) to propagate at different speeds Eardley:1973zuo . Furthermore, within a torsionless framework of metric-affine theory, up to eight independent polarization modes are possible Dong:2025ddi ; Dong:2025pyz , including additional shear-$x$ and shear-$y$ modes. The first multi-messenger observation, a binary neutron star coalescence GW170817 LIGOScientific:2017vwq and the associated gamma-ray burst GRB170817A Goldstein:2017mmi , places a stringent constraint on the speed of tensor modes $c_{t}$: $3\times 10^{-15}\leq c_{t}-1\leq 7\times 10^{-16}$ Savchenko:2017ffs . Clearly, its potential deviation from the speed of light is exceedingly minute. Considering the expansion of the universe and potential emission delays, it is now generally accepted that the propagation speed of tensor modes is equal to the speed of light. Numerous studies have examined GWs from an observational perspective, including imposing experimental constraints on theoretical models Baker:2017hug ; Creminelli:2017sry ; Dong:2022cvf , investigating their properties Liang:2022hxd ; Lai:2024fza ; Dong:2024zal ; Fan:2024pex ; Wang:2025hla ; Dong:2025ddi , and others Amarilo:2023wpn ; Jia:2024pdk ; DeFelice:2025ykh .
@@
In this paper, we focus on the cosmological perturbations for the Bumblebee model with a perfect fluid. First, using the background equations and the condition of an accelerated expanding universe, we perform a preliminary analysis of the parameter space and provide a brief discussion on dark energy. Then, we consider the scalar perturbations, vector perturbations, and tensor perturbations separately. We begin by deriving the second-order perturbation actions. Subsequently, based on the invariance of the linearized theory under infinitesimal coordinate transformations, we choose appropriate gauge conditions. We then eliminate the non-dynamical variables in Fourier space, which yields an effective action for the dynamical variables only. By taking the small-scale limit, we derive the conditions for the absence of ghost instability, Laplacian instability, and tachyonic instability, which further constrain the parameter space. Finally, within the constrained parameter space, we analyze the polarization modes, the number of degrees of freedom, the propagation speeds, and the amplitude relations among the different GW modes.
@@
This paper is structured as follows. In Sec. II, we briefly introduce the Bumblebee model and the Schutz-Sorkin action for a perfect fluid. In Sec. III, we perform a scalar-vector-tensor decomposition of the variables, analyze the background field equations, and discuss dark energy. In Sec. IV, we consider the tensor perturbations. We derive the second-order effective action, analyze the stability of the tensor perturbations, and constrain the Lorentz-violating parameter through GW observations. In Sec. V, we derive the second-order effective action for the vector perturbations and constrain the parameter space based on the stability conditions. In Sec. VI, we focus on the scalar perturbations. For three different gauge choices, we derive the stability conditions and constrain the parameter space. In Sec. VII, we analyze the polarization modes, the propagation speeds, and the amplitude relations of GWs within the constrained parameter space. Finally, we present our conclusion in Sec. VIII.
@@ II The Bumblebee model with a perfect fluid
The Bumblebee model Kostelecky:2003fs is a Lorentz-violating gravity theory that has a very simple form but encompasses interesting features, including rotation, boost, and CPT violations. Owing to its unique implications for our understanding of the universe, this model has been the subject of extensive research in GW physics and cosmology. It has been widely studied in the contexts of GWs Liang:2022hxd , black holes Liu:2024axg ; Filho:2022yrk , and cosmology vandeBruck:2025aaa .
The action of the Bumblebee model is given by Kostelecky:2003fs
$S=\int d^{4}x\left[\frac{1}{2\kappa}(R-2\Lambda+\xi B^{\mu}B^{\nu}R_{\mu\nu})-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})\right]+S_{m}.$ (1)
Here, $\kappa=8\pi G$, $\xi$ is a real coupling constant, $b^{2}$ is a real positive constant, and $B_{\mu\nu}=\nabla_{\mu}B_{\nu}-\nabla_{\nu}B_{\mu}$ is the field strength of the Bumblebee field $B_{\mu}$. The sign in $V(B^{\mu}B_{\mu}\pm b^{2})$ depends on whether $B_{\mu}$ is timelike or spacelike. The Bumblebee field gives rise to Lorentz violation through the potential $V$, which provides a nonzero vacuum expectation value $b_{\mu}$ for $B_{\mu}$. In this paper, we consider a spatially isotropic cosmological background. Thus, we take $b_{\mu}=(b_{t},0,0,0)$ with $b_{t}^{2}=b^{2}$.
@@
In this paper, we consider a perfect fluid that is minimally coupled to gravity and can be described by the Schutz-Sorkin action Schutz:1977df ; DeFelice:2009bx ; Bertolami:2008ab ; DeFelice:2016yws ; Kase:2018nwt
@@
Here, $\rho_{m}$ is the energy density, $n$ the particle number density, $J^{\mu}$ a vector density, and $\ell$ a scalar. The quantities $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$ are related to the intrinsic vector perturbations of the matter (see Refs. DeFelice:2009bx ; DeFelice:2016yws ).
@@
where $\rho_{m,n}=\partial\rho_{m}/\partial n$. One can show that the spatial components $U_{i}$ of $U_{\mu}$ can be decomposed into a scalar part and a divergence-free vector part. This decomposition holds even if $\rho_{m,n}$ is constant, which is consistent with Refs. DeFelice:2009bx ; Schutz:1970my . In the cosmological background, the divergence-free vector part of $U_{i}$ is generated by the scalars $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$.
@@ III Perturbations and cosmological background
$S$ $=$ $\int d^{4}x\left[\frac{1}{2\kappa}(R-2\Lambda+\xi B^{\mu}B^{\nu}R_{\mu\nu})-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})\right]$ (7)
$-\int d^{4}x\left[\sqrt{-g}\rho_{m}(n)+J^{\mu}(\partial_{\mu}\ell+\mathcal{A}_{1}\partial_{\mu}\mathcal{B}_{1}+\mathcal{A}_{2}\partial_{\mu}\mathcal{B}_{2})\right].$
Observations have shown that the current universe is very close to a spatially flat geometry WMAP:2006bqn . Therefore, we will analyze the equations of motion for the Bumblebee model within a spatially flat cosmological background
@@ III.1 Perturbations in a cosmological background
For a spatially homogeneous and isotropic universe, field perturbations can always be decomposed into spatial tensor, vector, and scalar parts, according to the transformation properties under 3-dimensional spatial rotation. This decomposition has been introduced in Refs. Flanagan:2005yc ; Jackiw:2003pm . Using this, the metric $g_{\mu\nu}$, the Bumblebee field $B_{\mu}$, the vector density $J^{\mu}$, and the scalar $\ell$ in a cosmological background can be decomposed as
@@
For $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$, we make the simplest choice that contains all the necessary information about the vector perturbations of matter DeFelice:2009bx ; DeFelice:2016yws

### v3

@@ Abstract
We constrain the parameter space of the Bumblebee model in a cosmological background and then investigate the properties of gravitational waves within the constrained parameter space. Our analysis reveals seven perturbative degrees of freedom in the cosmological background: two tensor, two vector, and two scalar modes, along with an additional mode from the matter sector. The stability conditions for all these modes are derived. By incorporating the observed accelerated expansion of the universe and the observational constraints on tensor gravitational waves, we derive bounds on the parameter space of the Bumblebee model. Our results indicate that the non-minimal coupling parameter $\xi$ must be non-positive, a constant background value $b_{t}$ of the Bumblebee field implies $\sigma\neq-\tfrac{1}{2}\xi$, and the Lorentz-violating parameter $\xi b^{2}$ has a lower bound on the order of $10^{-15}$. We then investigate the propagation characteristics and polarization modes of gravitational waves in both the small-scale and Minkowski limits. The propagation modes of gravitational waves in the Bumblebee model consist of two tensor modes, two vector modes, and one scalar mode. Notably, the tensor modes travel at subluminal speeds, whereas the vector and scalar modes propagate at superluminal speeds, when $\xi b_{t}^{2}\neq 0$. These results provide a concrete theoretical framework and specific observational signatures for testing Lorentz invariance in the gravitational sector with future gravitational-wave detectors.
@@ I Introduction
Since its proposal, general relativity (GR) has undergone rigorous tests in both weak-field and strong-field regimes. In the weak-field limit, its predictions have been precisely verified through observations and experiments, such as the precession of Mercury’s perihelion Einstein (1915), the deflection of light Dyson et al. (1923), and the Pound-Rebka experiment Pound and Rebka (1959, 1960). In the strong-field limit, the orbital decay of the Hulse-Taylor pulsar shows excellent agreement with the gravitational wave (GW) radiation predicted by GR Hulse and Taylor (1975); Taylor et al. (1979). The first direct detection of GWs, GW150914, confirms a key prediction of GR Abbott and others (2016b, a). Imaging of the black hole shadows of M87* and Sagittarius A* by the Event Horizon Telescope provides further support for the black hole solutions predicted by GR Akiyama and others (2019, 2022a, 2022b). This series of observations and experiments has tested the validity of GR across both the weak-field regimes and the highly non-linear strong-field regimes.
@@
Although GR has successfully withstood many experimental tests, several critical issues remain difficult to explain within its framework, such as the dark matter problem Smith (1936); Zwicky (1937), the dark energy problem Peebles and Ratra (2003), the problem of quantization ’t Hooft and Veltman (1974); Goroff and Sagnotti (1986), and the hierarchy problem Arkani-Hamed et al. (1998); Randall and Sundrum (1999a, b). This has led to questions within the physics community about whether GR is the ultimate theory of gravity, which has spurred research into modified gravity theories.
@@
The main approaches for constructing modified gravity theories include Nojiri and Odintsov (2011); Clifton et al. (2012):
  • (1)
    adding new fields: Brans-Dicke theory Brans and Dicke (1961), Horndeski theory Horndeski (1974), Bumblebee theory Kostelecky (2004), and so on Bekenstein (2004); Jacobson (2007); Heisenberg (2014); Geng and Lu (2016);
  • (2)
    considering higher-order derivatives: $f(R)$ theory Sotiriou and Faraoni (2010), $f(\mathcal{G})$ theory De Felice and Tsujikawa (2009), and so on Horava (2009b); Helpin and Volkov (2020);
  • (3)
    exploring higher dimensions: Kaluza-Klein theory Kaluza (1921), Randall-Sundrum theory Randall and Sundrum (1999a, b), and so on Rubakov and Shaposhnikov (1983); Arkani-Hamed et al. (1998);
  • (4)
    modifying geometry: Palatini-$f(R)$ theory Buchdahl (1970), $f(Q)$ theory Heisenberg (2024), and so on Hehl et al. (1995); Lu et al. (2020); Beltrán Jiménez et al. (2019);
  • (5)
    others: non-local theory Horava (2009a); Deser and Woodard (2007); Deffayet and Woodard (2009), spatially covariant gravity Gleyzes et al. (2013); Gao (2014b, a).
@@
Many of these modified gravity theories offer unique interpretations of GWs and the universe that differ from the interpretations provided by GR. For example, in general metric theories, up to six independent polarization modes of GWs are possible, significantly exceeding the two tensor modes predicted by GR Eardley et al. (1973). Certain modified gravity theories can effectively describe the evolution of the universe and provide a theoretical basis for explaining dark energy Nojiri and Odintsov (2006); Nojiri et al. (2017). Some modified gravity theories fit galaxy rotation curves well, offering viable alternatives to the dark matter hypothesis Famaey and McGaugh (2012); Capozziello and De Laurentis (2012). Therefore, theoretical consistency analyses and experimental tests are crucial for identifying a more complete description of gravity.
@@
A theoretically consistent modified gravity theory must be stable. A critical issue in this regard is the ghost problem. If ghost degrees of freedom, i.e., field modes carrying negative energy, are present, the theory will inevitably lead to vacuum instability or even catastrophic decay Woodard (2007). Thus, the ghost-free condition serves as an important tool for selecting modified gravity theories and constraining theoretical parameters De Felice et al. (2016); Kase and Tsujikawa (2019); Clough et al. (2022). Especially, the ghost problem is almost ubiquitous in higher-order modified gravity theories. The well-known Ostrogradsky theorem Ostrogradsky (1850) states that when a Lagrangian contains higher-order (second order or higher) time derivatives of dynamical variables, its Hamiltonian is usually bounded neither from above nor from below Woodard (2007). This implies that positive and negative parts could be excited to arbitrarily high energies, leading to vacuum instability Woodard (2007). There are also other studies concerning theoretical consistency, such as the initial value problem and the stability of evolution Kreiss and Lorenz (1989); Papallo and Reall (2017); Bluhm et al. (2008); Clough et al. (2022).
@@
For the experimental tests of modified gravity theories, one of the most direct and promising approaches is the detection and characterization of GW signals. The polarization modes and propagation speeds of GWs are two critically important characteristics that typically vary across different modified gravity theories, making them direct probes for testing gravity theories. In GR, GWs propagate at the speed of light and exhibit only two tensor polarization modes: the plus mode ($P_{+}$) and the cross mode ($P_{\times}$). However, a general metric theory in four-dimensional spacetime allows for GWs up to six independent polarization modes (the additional ones being breathing ($P_{b}$), longitudinal ($P_{l}$), vector-$x$ ($P_{x}$), and vector-$y$ ($P_{y}$) modes) to propagate at different speeds Eardley et al. (1973). Furthermore, within a torsionless framework of metric-affine theory, up to eight independent polarization modes are possible Dong et al. (2025a, b), including additional shear-$x$ and shear-$y$ modes. The first multi-messenger observation, a binary neutron star coalescence GW170817 Abbott and others (2017b) and the associated gamma-ray burst GRB170817A Goldstein and others (2017), places a stringent constraint on the speed of tensor modes $c_{t}$: $3\times 10^{-15}\leq c_{t}-1\leq 7\times 10^{-16}$ Savchenko and others (2017). Clearly, its potential deviation from the speed of light is exceedingly minute. Considering the expansion of the universe and potential emission delays, it is now generally accepted that the propagation speed of tensor modes is equal to the speed of light. Numerous studies have examined GWs from an observational perspective, including imposing experimental constraints on theoretical models Baker et al. (2017); Creminelli and Vernizzi (2017); Dong et al. (2023), investigating their properties Liang et al. (2022); Lai et al. (2024); Dong et al. (2025c); Fan et al. (2025); Wang et al. (2025); Dong et al. (2025a), and others Amarilo et al. (2024); Jia et al. (2024); De Felice and Hell (2025).
@@
In this paper, we focus on the cosmological perturbations for the Bumblebee model with a perfect fluid. First, using the background equations and the condition of an accelerated expanding universe, we perform a preliminary analysis of the parameter space and provide a brief discussion on dark energy. Then, we consider the scalar perturbations, vector perturbations, and tensor perturbations separately. We begin by deriving the second-order perturbation actions. Subsequently, based on the invariance of the linearized theory under infinitesimal coordinate transformations, we choose appropriate gauge conditions. We then eliminate the non-dynamical variables in Fourier space, which yields an effective action for the dynamical variables only. Starting from the effective action, we first derive the ghost-free conditions for the perturbations and then, in the small-scale limit, analyze the stability and propagation characteristics of GWs, which further constrain the parameter space. Finally, within the constrained parameter space, we analyze the polarization modes, the number of degrees of freedom, the propagation speeds, and the amplitude relations among the different GW modes.
@@
This paper is structured as follows. In Sec. II, we briefly introduce the Bumblebee model and the Schutz-Sorkin action for a perfect fluid. In Sec. III, we perform a scalar-vector-tensor decomposition of the variables, analyze the background field equations, and discuss dark energy. In Sec. IV, we consider the tensor perturbations. We derive the second-order effective action, analyze the stability of the tensor perturbations, and constrain the Lorentz-violating parameter through GW observations. In Sec. V, we derive the second-order effective action for the vector perturbations and constrain the parameter space based on the stability conditions. In Sec. VI, we focus on the scalar perturbations. For three different gauge choices, we derive the stability conditions, analyze the propagation characteristics of GWs, and constrain the parameter space. In Sec. VII, we analyze the polarization modes, the propagation speeds, and the amplitude relations of GWs within the constrained parameter space. Finally, we present our conclusion in Sec. VIII.
@@ II The Bumblebee model with a perfect fluid
The Bumblebee model Kostelecky (2004) is a Lorentz-violating gravity theory that has a very simple form but encompasses interesting features, including rotation, boost, and CPT violations. Owing to its unique implications for our understanding of the universe, this model has been the subject of extensive research in GW physics and cosmology. It has been widely studied in the contexts of GWs Liang et al. (2022), black holes Liu et al. (2025); Filho et al. (2023), and cosmology van de Bruck et al. (2025).
The action of the Bumblebee model is given by Kostelecky (2004)
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa}(R-2\Lambda+\xi B^{\mu}B^{\nu}R_{\mu\nu}+\sigma B^{\mu}B_{\mu}R)-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})\right]+S_{m}.$ (1)
Here, $\kappa=8\pi G$, $b^{2}$ is a real positive constant, $B_{\mu\nu}=\nabla_{\mu}B_{\nu}-\nabla_{\nu}B_{\mu}$ is the field strength of the Bumblebee field $B_{\mu}$, and $\xi$ and $\sigma$ are real coupling constants. The sign in $V(B^{\mu}B_{\mu}\pm b^{2})$ depends on whether $B_{\mu}$ is timelike or spacelike. The Bumblebee field gives rise to Lorentz violation through the potential $V$, which provides a nonzero vacuum expectation value $b_{\mu}$ for $B_{\mu}$. In this paper, we consider a spatially isotropic cosmological background. Thus, we take $b_{\mu}=(b_{t},0,0,0)$ with $b_{t}^{2}=b^{2}$.
@@
In this paper, we consider a perfect fluid that is minimally coupled to gravity and can be described by the Schutz-Sorkin action Schutz and Sorkin (1977); De Felice et al. (2010); Bertolami et al. (2008); De Felice et al. (2016); Kase and Tsujikawa (2018)
@@
Here, $\rho_{m}$ is the energy density, $n$ the particle number density, $J^{\mu}$ a vector density, and $\ell$ a scalar. The quantities $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$ are related to the intrinsic vector perturbations of the matter (see Refs. De Felice et al. (2010, 2016)).
@@
where $\rho_{m,n}=\partial\rho_{m}/\partial n$. One can show that the spatial components $U_{i}$ of $U_{\mu}$ can be decomposed into a scalar part and a divergence-free vector part. This decomposition holds even if $\rho_{m,n}$ is constant, which is consistent with Refs. De Felice et al. (2010); Schutz (1970). In the cosmological background, the divergence-free vector part of $U_{i}$ is generated by the scalars $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$.
@@ III Perturbations and cosmological background
$S$ $=$ $\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa}\left(R-2\Lambda+\xi B^{\mu}B^{\nu}R_{\mu\nu}+\sigma B^{\mu}B_{\mu}R\right)-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})\right]$ (7)
$-\int d^{4}x\left[\sqrt{-g}\rho_{m}(n)+J^{\mu}\left(\partial_{\mu}\ell+\mathcal{A}_{1}\partial_{\mu}\mathcal{B}_{1}+\mathcal{A}_{2}\partial_{\mu}\mathcal{B}_{2}\right)\right].$
Observations have shown that the current universe is very close to a spatially flat geometry Spergel and others (2007). Therefore, we will analyze the equations of motion for the Bumblebee model within a spatially flat cosmological background
@@ III.1 Perturbations in a cosmological background
For a spatially homogeneous and isotropic universe, field perturbations can always be decomposed into spatial tensor, vector, and scalar parts, according to the transformation properties under 3-dimensional spatial rotation. This decomposition has been introduced in Refs. Flanagan and Hughes (2005); Jackiw and Pi (2003). Using this, the metric $g_{\mu\nu}$, the Bumblebee field $B_{\mu}$, the vector density $J^{\mu}$, and the scalar $\ell$ in a cosmological background can be decomposed as
@@
For $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$, we make the simplest choice that contains all the necessary information about the vector perturbations of matter De Felice et al. (2010, 2016)